Vicky's mother is 30 years older than she is. In 12 years' time, Vicky's mother will be 3 times as old as her. How old will Vicky be in 5 years' time?
Answer by dkppathak(439)

: Vicky's mother is 30 years older than she is. In 12 years' time, Vicky's mother will be 3 times as old as her. How old will Vicky be in 5 years' time?
0 solutions
let vicky age is x yrs
mother age is x+30
12 yrs from now
viky age will be x+12 and mother will be x+30+12=x+42
x+42=3( x+12)
x+42=3x+36
42-36=2x
x=6/2=3
present age of viky is 3 yrs and 5 yrs now will be 5+3=8 yrs
answer 8 yrs